# Transfer ABCD Matrix for Time-Varying Media and Time Crystals

### v1

@@ IV Conclusion
In this paper, we have introduced a formal definition of the transfer ABCD parameters applied to two-port time-varying systems and time crystals. The definition comes from the rearrangement of the fields $\mathbf{D}$ and $\mathbf{B}$ at the input and output, so that the temporal boundary conditions are fulfilled. Thus, the governing temporal transfer matrix [eq. (2)] can be cascaded in a similar fashion to conventional ABCD parameters. Then, we have computed the ABCD parameters of a temporal transmission line (temporal slab) in eq. (9), and its associated scattering parameters in eqs. (11)-(14). Moreover, we have discussed on the continuity of the momentum density vector $\mathbf{D}\times\mathbf{B}$, and derived an analytical expression [eq. (22)] to compute the dispersion diagram ($\omega$ vs $k$) in time crystals via the ABCD formalism. Finally, we have tested the theory by computing the dispersion diagram of a time crystal and the scattering parameters of a temporal multilayer structure. The results presented in Section III validate the use of the proposed temporal transfer ABCD matrix and open up new possibilities in the analysis and design of time-varying system.

### v2

@@ IV Conclusion
In this paper, we explore a formal definition of the transfer ABCD parameters applied to two-port time-varying systems and time crystals. The definition comes from microwave theory concepts and the rearrangement of the fields $\mathbf{D}$ and $\mathbf{B}$ at the input and output, so that the temporal boundary conditions are fulfilled. Thus, the governing temporal transfer matrix [eq. (2)] can be cascaded in a similar fashion to conventional ABCD parameters. Then, we have computed the ABCD parameters of a temporal transmission line (temporal slab) in eq. (9), and its associated scattering parameters in eqs. (11)-(14). Moreover, we have discussed on the continuity of the momentum density vector $\mathbf{D}\times\mathbf{B}$, and derived an analytical expression [eq. (22)] to compute the dispersion diagram ($\omega$ vs $k$) in time crystals via the ABCD formalism. Finally, we have tested the theory by computing the dispersion diagram of a time crystal and the scattering parameters of a temporal multilayer structure. The results presented in Section III validate the use of the proposed temporal transfer ABCD matrix and open up new possibilities in the analysis and design of time-varying systems.
Acknowledgments
This work was supported by grant PID2020-112545RB-C54 funded by MCIN/AEI/10.13039/501100011033 and by European Union NextGenerationEU/PRTR. It has also been supported by TED2021-131699B-I00, PDC2023-145862-I00, and IJC2020- 043599-I, funded by MCIN/AEI/10.13039/501100011033 and by The European Union NextGenerationEU/PRTR. It has also been supported in part by Consejería de Universidad, Investigación e Innovación of Junta de Andalucía through grant EMERGIA 23-00235”.